OpenStudy (anonymous):

Find the vector, parametric and scalar equation of the plane passing through A(3, 5, 2), B(0, 5, −1) and C(1, 5, −3)

OpenStudy (anonymous):

Find your range and then use your kernel

OpenStudy (amistre64):

the vector equation is just "normal" dotted to one of the given points

OpenStudy (amistre64):

i cant say that im familiar with the other two forms .....

OpenStudy (anonymous):

Amistre are you like 22? :D

OpenStudy (amistre64):

pfft, i havent been 22 in ages :)

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

So I cross two points then dot it with the third?

OpenStudy (amistre64):

cross 2 vectors to define the normal vector

OpenStudy (amistre64):

you are given 3 points to define vectors with

OpenStudy (amistre64):

the ax+by+cz +d = 0 is the scalar format

OpenStudy (amistre64):

would parametric be the matrix form?

OpenStudy (anonymous):

So it'd be (AxB) . C?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Parametric is when you have its in rows. Yeah somewhat like Matrix form. I know how to get the respect forms I'm just not sure what to do to get started

OpenStudy (amistre64):

A cross B, dotted to C

OpenStudy (amistre64):

A(3, 5, 2), B(0, 5, −1) C(1, 5, −3) -5 +1 -5 +1 -5 +1 ---------------------------------- 3 0 3 0 0 0 1 0 -2

OpenStudy (anonymous):

Parametric is just the vector form broken up

OpenStudy (amistre64):

x 3 1 y 0 0 z 3 -2 x = 0 y = -(6-3) z = 0

OpenStudy (anonymous):

When I got (AxB) . C I end up with -45

OpenStudy (amistre64):

C is a generic vector (x-xo, y-yo, z-zo)

OpenStudy (anonymous):

So after I got this -45 how do I get the forms?

OpenStudy (amistre64):

um, you dont get -45 to start with ...

OpenStudy (anonymous):

Well A cross B is (-15,3,15) dot (1,5,-3) is -45

OpenStudy (amistre64):

C = ( \(x-1,y-5,z+3\)~)

OpenStudy (amistre64):

your operations are not correct

OpenStudy (anonymous):

Okay lemme try that.

OpenStudy (anonymous):

Nvm could you please explain

OpenStudy (amistre64):

your points given all have y=5; which means that this plane is parallel to y=5

OpenStudy (amistre64):

the equation is: x=x, z=z, y=5

OpenStudy (amistre64):

A(3, 5, 2) B(0, 5, −1) C(1, 5, −3) ^^ all ys are the same this is a plane parallel to the zx plane

OpenStudy (anonymous):

Okay. So how do I go about getting the vector equation?

OpenStudy (amistre64):

your normal is just the (0,1,0) vector dotted to the generic C

OpenStudy (anonymous):

When you dot something isn't it a scalar value? How do you dot something with an x,y,z variable in it?

OpenStudy (anonymous):

Wouldn't I just end up with y-5?

OpenStudy (amistre64):

take generic point (x,y,z), and create a vector to if from C(1,5,-3) this produces all possible vectors from C to every single point (x-1,y-5,z+3) defines all vectors from C to a given general point

OpenStudy (amistre64):

the dotproduct to a normal defines all vectors from C that are perpendicular (orthogonal) to the normal vector

OpenStudy (amistre64):

a plane is defined as all the vectors that are perpendicular to a given vector

OpenStudy (amistre64):

the vector equation is : \(\vec n\cdot \vec C\)\[(0,1,0)\cdot (x-1,y-5,z+3)\]

OpenStudy (amistre64):

the parametric is the result of that \[0(x-1)+1(y-5)+0(z+3)=0\]

OpenStudy (amistre64):

or is that the scalar?

OpenStudy (amistre64):

the parametric might be: x = 0 + 1x + 0z y = 5 + 0x + 0z z = 0 + 0x + 1z

OpenStudy (anonymous):

Umm this is what I did. Please let me know if it makes sense. a(x-xo) + b(y-y0) + c(z-z0) vector 1 = AB = (-3,0,-3) vector 2 = AC = (-2,0,-5) v1 x v2 = (0,-9,0) 0(x-1) -9(y-5) + 0(z+3) = 0 -9y-45 = 0 -9y = 45 y = -5 And that would be the equation of the plane. Does that make sense?

OpenStudy (amistre64):

it almost makes sense ... y = 5 should be the outcome so you most likely ignored a negative along the way

OpenStudy (amistre64):

-9(-5) = +45

OpenStudy (anonymous):

Ahh. Yeah. Makes sense now.